A 45-kg sample of water absorbs 345 kJ of heat. If the water was initially at 22.1 °C, what is its final temperature?
1
Expert's answer
2020-09-07T08:20:14-0400

From the formula for finding heat

Q=c×m×∆T:

∆T=Q/c×m.

c - heat capacity of water is known from the table data:

c=4182J/kg°C.

Units of heat should be turned into Joules:

Q=345×1000=345000(J)

∆T=345000/45×4182=1.8°.

If the heat is absorbed, the final temperature is higher than the initial one:

T=T0+∆T=22.1°+1.8°=23.9°.